OpenStudy (anonymous):

The perimeter of a rectangle is 50cm and the length is 4cm longer than the width. Write and solve system of equations using the substituution method. a) width = 10cm, length=14cm b) width = 10 1/2 cm, length = 14 1/2cm c) width = 11cm, length = 15cm d) width = 23cm, length = 27cm

OpenStudy (nubeer):

width = x length =x+4 perimeter = 50 do you know how to find the permiter of rectangle?

OpenStudy (anonymous):

@nubeer not really lol

OpenStudy (nubeer):

permieter of rectangle = 2(length +width)

OpenStudy (nubeer):

now plug in values to find the value of x

OpenStudy (anonymous):

@nubeer so 2(2x+4)

OpenStudy (nubeer):

yes = 50

OpenStudy (anonymous):

@nubeer so i dont understand how this helps me find the perimeter

OpenStudy (anonymous):

@nubeer i meant length and width

OpenStudy (nubeer):

you dont have to find perimeter.. perimeter is already given u have to find the value of x first.

OpenStudy (nubeer):

ohh well 2(2x+4)=50 4x+8 =50 can you find x now?

OpenStudy (nubeer):

when u find the value of x.. you know width =x so x would be your width

OpenStudy (anonymous):

@nubeer so the answer is b?

OpenStudy (nubeer):

yes it seems so :)

OpenStudy (anonymous):

@nubeer thank you!

OpenStudy (nubeer):

you are welcome